OpenStudy (anonymous):

Deandre is estimating the weight of his mountain bike. The actual weight of his mountain bike is 15.5kg . Deandre's estimate is 13kg . Find the absolute error and the percent error of Deandre's estimate. If necessary, round your answers to the nearest tenth.

OpenStudy (anonymous):

The actual weight of his mountain bike is 14.5 NOT 15.5 , sorry

OpenStudy (ghostgate):

Its okay.

OpenStudy (ghostgate):

Hmmm, I can't really seem to help. Is there a formula to mesure things without having comparisons or any other number?

OpenStudy (anonymous):

nope nothing

OpenStudy (anonymous):

they only give the question

OpenStudy (ghostgate):

Wait a second.

OpenStudy (anonymous):

Absolute: 14.5 - 13 = 1.5 kg Percentage: 1.5 / 14.5 = 3 / 29

OpenStudy (ghostgate):

Okay, so his bike actually weighs 14.5 kg and Dandre thought it weighed 13 kg. So he was 1.5 kg off. However he was 10.4% off. In order to find the percentage difference. Take 100 divide it by 14.5 and multiply by 1.5 to get your answer. Anyways he is 10.4% off.

OpenStudy (anonymous):

Percentage = 3 / 29 = 10.344%

OpenStudy (ghostgate):

So as ColossalMagnitude said. The absolute number he was off by was 1.5 which you do by taking 14.5 - 13 = 1.5.

OpenStudy (anonymous):

so percent error is 10.4

OpenStudy (anonymous):

Good job, Ghostgate. I think you should get the medal because you explained it really well, even though I said the answer first.

OpenStudy (anonymous):

and absolute error is 1.5

OpenStudy (ghostgate):

Thanks ColossalMagnitude! So the Percent error is 10.4% due to how you take 100 / 14.5 x 1.5(Since 1.5 is the difference) and you get 10.4%. The absolute error which is the difference between the two is 1.5. You get this by taking 14.5 - 13 = 1.5 since 1.5 is the amount of numbers inbetween the two numbers.

OpenStudy (anonymous):

incorrect

OpenStudy (anonymous):

percent error is not 10.4

OpenStudy (anonymous):

says one of the digits are wrong

OpenStudy (ghostgate):

Oh wait its 10.3. I'm sorry. It was like 10.33 something so it would be 10.3% Sorry man. Can you redo it?

OpenStudy (anonymous):

We said it was 10.3 slimjim -.- or atleast I did.

OpenStudy (anonymous):

yea its fine i submitted again

OpenStudy (ghostgate):

Ranga, in percentages you don't have negatives and positive percentages.

OpenStudy (anonymous):

yea sorry i thought it gets rounded to 10,4 like ghost said

OpenStudy (ghostgate):

Well it could of been 10.4 because the answer is like 10.3448___etc. So you would round to 10.345 to 10.35 to 10.4, but they didn't imply that.

OpenStudy (anonymous):

nothing bad happened , Thanks both of you

OpenStudy (ghostgate):

You're welcome! Have a G'day!

OpenStudy (ghostgate):

Ah... I stand corrected, thanks for pointing that out Ranga. I should give you a medal! No jk. I can't. :P

OpenStudy (ranga):

np.

OpenStudy (ghostgate):

If there's any other questions slimjim77. Feel free to ask by messaging one of us okay?

OpenStudy (anonymous):

i will , thanks